OpenStudy (anonymous):

Last question, 2 parts: Part 1: For the following system of equations, write your own real world scenario that describes what is happening. Use complete sentences and correct grammar in your scenario. Part 2: Solve the system and explain what the results mean according to your scenario. 2x + y = 10 3x + 4y = 25

Parth (parthkohli):

Systems never work in real life :S they just are to add pressure to the student-life

Parth (parthkohli):

If x is the number of apples and y is the number of oranges: 2 times the apples you have and the number of apples you have would add and make 10 fruits. Also, three times the apples you have and 4 times the oranges you have would add and make 25 fruits.

Parth (parthkohli):

^ That's the best I could think of :P

OpenStudy (anonymous):

Its good :) I like it .

OpenStudy (anonymous):

the equation ares 2x+y=10......1 3x+4y=25.....2 from equation 1 x=10-y/2 put it in 2nd equ 3(10-y/2)+4y=25 30-3y/2+4=25 now by LCM 30-3y+8y/2=25 by cross multiply 30+5y=50 5y=20 y=4 now put y=4 in equ 1 to get value of x 2x+4=10 2x=10-4 2x=6 x=3 if youu like my answer please give me medal. Thanks!

OpenStudy (anonymous):

is this part 2?

OpenStudy (anonymous):

yes

Parth (parthkohli):

Oops.. asking for a medal is the worst thing you could ever do. A tip from me.

Parth (parthkohli):

I supposed that you could solve that.

OpenStudy (anonymous):

@AmberLynn02 did you like my answer?